# Data-Driven Prescriptive Analytics Applications: A Comprehensive Survey

### v1

@@ X-C Predict-Then-Prescribe (PTP) workflows
Aside from certain advanced extensions falling under the PWP pattern, Linear Programming generally implies a PTP workflow in the set of surveyed papers. The technical reason behind this phenomenon might be that popular, useful forecasting models, like tree-based ensembles or Logistic Regression for instance (cf. Section IX-A) tend to have non-linear, black-box, and/or non-parametric properties that are incompatible with the restrictions of linear modelling. The only sure-fire way to combine such forecasting models with Linear Programming is therefore to pre-compute forecasts as linear optimization model coefficients.
@@
After having obtained exemplars [82, 53, 47], scheduling requirements [109, 75], found defects to be remediated [111, 112, 125], possible yields from new employees or loans [145, 78], or potential uplift per customer [59, 60, 58, 61, 62] from predictive technologies like DM/ML, heuristics and problem domain reasoning can be applied to provide prescriptions in a subsequent phase (cf. Section IX-D), analogously to the Predict-Then-Optimize pattern. As it turns out, all surveyed papers relying solely on EXP for prescriptions utilize PTP workflows. This could perhaps be explained by the notion that implementing sound domain reasoning and heuristics with respect to complex forecast models generally is easier, if the input for this analysis is simply the forecast value as is.
As for prescriptive DM/ML, prescriptions through Nearest Neighbour Queries [108, 77], feature selection algorithms [114], or trained ML models [67, 66, 128, 152] can be obtained similarly to the EXP case, after upstream predictive models have provided their input (cf. Section IX-A). As we shall see shortly, prescriptive DM/ML however also comes in different shapes within the PWP pattern.
X-D Predicting-While-Prescribing (PWP) workflows
@@
The need for online predictions from what is usually predictive DM/ML models (cf. Figure 7), somewhat limits the array of applicable optimization methods able to cope with these cases. As found in this survey, Heuristic Optimization methods [86, 106, 139, 116, 144, 126, 127, 117, 48, 140] were strongly associated with the PWP pattern, yet Exhaustive Search was also utilized [92, 57, 88, 143], and one paper used online gradient estimations of a Support Vector Regression model to solve a convex optimization problem [83] (cf. Section IX-B).
@@
Analogously to PWP solutions using prescriptive OPT, SIM solutions using Simulation-Based Optimization (cf. Section IX-E) also exclusively follow a PWP pattern [135, 54, 74, 110, 71, 149].
As seen in Figure 9, remaining PWP workflows all rely on some form of DM/ML. Similarly to the aforementioned cases, papers using techniques from eXplainable AI to form prescriptions using SHAP values or DiCE [148, 50] (cf. Section IX-A) need to run online predictions to obtain the required output from the aforementioned methods, in that these methods are designed to work for a variety of black-box models in need of explainability [163, 164].
In a sense, Reinforcement Learning embodies the PWP pattern. Some reinforcement learning algorithms solve predictive and prescriptive problems concurrently - the value and policy function, respectively estimating the prospect of being in a state and the optimal way to navigate from it, being ”moving targets” for each other until convergence [165, p. 97]. Whether they use data-driven reinforcement learning as a stand-alone approach [55, 65, 64, 113, 51] or supplement it with additional predictive methods used in an online fashion [122, 120, 68], papers using reinforcement learning all at least embody the PWP pattern in such a sense.
@@
To give a concrete example, one paper trained the coefficients for their prescriptive model, which calculates Random Forest kernelized weights for the input vector with respect to a set of historical data points (cf. Section IX-A), by solving a convex optimization problem involving said Random Forest kernel [79]. In another paper, it was proposed to solve several instances of the core optimization problem of the application at hand, and then use this dataset of problem-solution pairs to train a tree-based ensemble to predict optimal solutions for new input vectors, which experiments indicated to be an viable approach compared to alternatives [80].
XI RQ2 synthesis
As Section IX and X have covered, several definite methodological trends characterize the state of current DPSA application research, both in terms of individual methods and their combinations.
DPSA applications can be said to pivot around five method types, or ”clusters” if such imagery is preferred, among which DM/ML and OPT dominate (cf. Figure 4), yet with several viable alternatives in the shape of PROB, EXP, and SIM. While a pronounced emphasis on tree-based ensembles for predictions and Linear Programming methods for prescriptions was found (cf. Section IX-A and IX-B), individual method choices varied widely between applications, exhibiting a textbook level of coverage for the most popular method types. To our knowledge, the various prescriptive EXP methods described in Section IX-D have not been covered explicitly elsewhere, despite the popularity of prescriptive EXP (cf. Figure 7).
In terms of how methods were combined, we found that DPSA applications pivot around a relatively small number of method combinations among the aforementioned types (cf. Section X-A), utilizing primarily DM/ML and secondarily PROB for predictions, and primary OPT, supplemented by DM/ML and EXP, for prescriptions. We observe from Figure 7 that there is seemingly a power law at play in terms of popularity of individual method types for prediction/prescription, as well as their combinations, with (combinations of) DM/ML and OPT clearly being the most popular, yet contested EXP for prescriptions and PROB for predictions.
As for workflow patterns, we identified two generic patterns for structuring DPSA workflows, dividing the set of surveyed papers into two evenly-sized parts (cf. Section X-B). The Predict-Then-Prescribe pattern is already well-known within the methodological PSA literature in some form [22], as a sequential pattern with a clear-cut separation between prediction and prescription. The alternative Predicting-While-Prescribing pattern, as introduced in this survey, breaks the mold of PTP, bringing prediction and prescription closer together by simultaneous relationships (iterative and/or overlapping), with recent methodological developments even unifying data-driven prediction and automatic prescription in one model.
@@ XI RQ2 synthesis
XII RQ3: Possible research directions
Based on pervasive patterns found in the set of surveyed papers, we derived four possible research directions (RD1-4) to address RQ3, presented in the following subsections.
XII-A The promise of Big DPSA
Big Data can be said to form a significant part of a narrative surrounding PSA, in which PSA, extending the reach of its descriptive and predictive counterparts (cf. Figure 1), is the third and final piece to unlock maximum Value from Big Data in a fully automated, large-scale analytics infrastructure [166, 27, 24]. A substantial number of surveyed papers correspondingly discuss the potentials of Big Data in their opening or closing sections [92, 135, 86, 59, 80, 109, 134, 113, 111, 79, 110]. Nonetheless, hardly any DPSA applications in this survey work with what we would even consider to be Big Data.
To clarify our own definition of the phenomenon, we prefer to avoid arbitrary distinctions counted in bytes or cardinalities and instead define data-intensive processing by the need for distributed processing - as exemplified by the primary use cases of distributed data ingestion, analytics and storage systems like Apache Kafka, Spark, and Cassandra [167, 35, 168]. In other words, if a dataset can feasibly be analyzed on one workstation/server with respect to application requirements, as done in most surveyed papers (e.g., [84, 106, 123, 126, 75, 81]), then any standard analytics solution like Excel and MySQL might suffice, and said dataset is not considered Big.
A few papers describe their own work as involving Big Data or evidently use distributed processing to handle large datasets [89, 121, 141, 77]. Yet these papers tend to omit implementation details, in some cases possibly for confidential reasons [141], while others only analyze single-node-sized datasets [77].
@@ XII-A The promise of Big DPSA
XII-B Under-explored application domains
As found in connection to RQ1 (cf. Section VIII), DPSA has already found use in a variety of identified application domains. Still, several prominent business domains have been left (almost) unexplored by the set of surveyed papers as a whole. Using UN’s International Standard Industrial Classification (ISIC), revision 4, [169], we have identified eleven business domains in which at most three DPSA applications were found in this survey. Leaving out the miscellaneous sections S, T, and U [169,  p. 43], the eight remaining business domains and section letters, along with associated surveyed papers, if any, are the following:
@@ XII-B Under-explored application domains
XII-C Methodological alternatives to (mixed-)integer linear programming
While covering OPT methods in Section IX-B, we found that Linear Programming, and in particular (Mixed-)Integer Linear Programming ((M)ILP) algorithms, were the most dominant class of methods by far. What we didn’t dwell on, for the sake of brevity, is the frequently recurring observation that one does not simply use (M)ILP for DPSA. Among DPSA applications there is a strong tendency of various difficulties arising whenever authors try to apply (M)ILP to real-world problems, calling for improvised modelling tricks, heuristics, and workflows.
@@ XII-C Methodological alternatives to (mixed-)integer linear programming
Another challenge with (M)ILP when considering DPSA specifically, is the popularity of predictive DM/ML (cf. Figure 7), in particular non-parametric or non-linear models (cf. Section IX-A). Limited options for embedding such models into linear optimization models makes (M)ILP less expressive, to the possible detriment of solution quality, e.g., as found in [140].
@@
While we recognize that it might be impossible to design a perfect alternative to current prescriptive methods, we propose as a possible research direction to:
RD3: Develop scalable, expressive, usable, and reliable methods for DPSA applications.
XII-D Flexible tools for DPSA
From our analyses concerning RQ2 (cf. Section IX and X), we note that DPSA applications display a high degree of methodological diversity, both in terms of individual methods and how they are combined. Accommodating such a wide variety of workflows poses a challenge to developers of dedicated DPSA tools: As is evident from the aforementioned sections, a PTP pattern with ML for predictions and Linear Programming for prescriptions is but one way to structure a DPSA workflow among many. As for (M)ILP specifically, we furthermore saw in Section XII-C that significant workflow variations even exist within this one category.
@@ XII-D Flexible tools for DPSA
XIII Conclusion
The primary objective of this survey was to investigate where the many applications of DPSA have been and where they are going (cf. Section I). In this endeavour, we decided to delimit our scope to Data-Driven PSA, a field deemed to have a sound basis in the historical origins of PSA as well as recent calls for papers (cf. Section IV).
In answering our three research questions in Section VIII-XII, we have provided new conceptual models of the field as a whole, both in terms of problem types (cf. Section VIII), method types (cf. Section IX), and workflow patterns (cf. Section X). Based on pervasive Gordian Knots found within the set of surveyed papers, we have furthermore provided a foundation for future work by proposing a number of possible research directions (cf. RD1-4 from Section XII), both in terms of new methods, tools, and use cases. We can mnemonically summarize the key contributions in our conceptual framework as 10 applications domains, 5 method types, 2 generic workflow patterns, and 4 research directions. These contributions form a comprehensive model of the current state of the DPSA application research field and where it might be going in the future.
Looking ahead, we believe that this survey can be a useful guide for new recruits and veterans alike. As found in Section XII, there are still ample opportunities for developing new methods, tools, and use cases for DPSA applications with a positive impact on the real world. Revisiting our opening quote in Section I, we hope that upcoming DPSA applications will continue to find new ways to create the future by predicting it.

### v2

@@ 10.3 Predict-Then-Prescribe (PTP) workflows
Aside from certain advanced extensions falling under the PWP pattern, Linear Programming generally implies a PTP workflow in the set of surveyed papers. The technical reason behind this phenomenon might be that popular, useful forecasting models, like tree-based ensembles or Logistic Regression for instance (cf. Section 9.1) tend to have non-linear, black-box, and/or non-parametric properties that are incompatible with the restrictions of linear modelling. The only sure-fire way to combine such forecasting models with Linear Programming is therefore to pre-compute forecasts as linear optimization model coefficients.
@@
After having obtained exemplars [85, 56, 54], scheduling requirements [112, 78], found defects to be remediated [114, 115, 128], possible yields from new employees or loans [148, 81], or potential uplift per customer [62, 63, 61, 64, 65] from predictive technologies like DM/ML, heuristics and problem domain reasoning can be applied to provide prescriptions in a subsequent phase (cf. Section 9.4), analogously to the Predict-Then-Optimize pattern. As it turns out, all surveyed papers relying solely on EXP for prescriptions utilize PTP workflows. This could perhaps be explained by the notion that implementing sound domain reasoning and heuristics with respect to complex forecast models generally is easier, if the input for this analysis is simply the forecast value as is.
As for prescriptive DM/ML, prescriptions through Nearest Neighbour Queries [111, 80], feature selection algorithms [117], or trained ML models [70, 69, 131, 155] can be obtained similarly to the EXP case, after upstream predictive models have provided their input (cf. Section 9.1). As we shall see shortly, prescriptive DM/ML however also comes in different shapes within the PWP pattern.
10.4 Predicting-While-Prescribing (PWP) workflows
@@
The need for online predictions from what is usually predictive DM/ML models (cf. Figure 9), somewhat limits the array of applicable optimization methods able to cope with these cases. As found in this survey, Heuristic Optimization methods [89, 109, 142, 119, 147, 129, 130, 120, 53, 143] were strongly associated with the PWP pattern, yet Exhaustive Search was also utilized [95, 60, 91, 146], and one paper used online gradient estimations of a Support Vector Regression model to solve a convex optimization problem [86] (cf. Section 9.2).
@@
Analogously to PWP solutions using prescriptive OPT, SIM solutions using Simulation-Based Optimization (cf. Section 9.5) also exclusively follow a PWP pattern [138, 57, 77, 113, 74, 152].
As seen in Figure 13, remaining PWP workflows all rely on some form of DM/ML. Similarly to the aforementioned cases, papers using techniques from eXplainable AI to form prescriptions using SHAP values or DiCE [151, 52] (cf. Section 9.1) need to run online predictions to obtain the required output from the aforementioned methods, in that these methods are designed to work for a variety of black-box models in need of explainability [166, 167].
In a sense, Reinforcement Learning embodies the PWP pattern. Some reinforcement learning algorithms solve predictive and prescriptive problems concurrently - the value and policy function, respectively estimating the prospect of being in a state and the optimal way to navigate from it, being ”moving targets” for each other until convergence [168, p. 97]. Whether they use data-driven reinforcement learning as a stand-alone approach [58, 68, 67, 116, 50] or supplement it with additional predictive methods used in an online fashion [125, 123, 71], papers using reinforcement learning all at least embody the PWP pattern in such a sense.
@@
To give a concrete example, one paper trained the coefficients for their prescriptive model, which calculates Random Forest kernelized weights for the input vector with respect to a set of historical data points (cf. Section 9.1), by solving a convex optimization problem involving said Random Forest kernel [82]. In another paper, it was proposed to solve several instances of the core optimization problem of the application at hand, and then use this dataset of problem-solution pairs to train a tree-based ensemble to predict optimal solutions for new input vectors, which experiments indicated to be an viable approach compared to alternatives [83].
11 Synthesis on methods
As Sections 9 and 10 have covered while addressing RQ2, several definite methodological trends characterize the state of current DPSA application research, both in terms of individual methods and their combinations.
DPSA applications can be said to pivot around five method types, or ”clusters” if such imagery is preferred, among which DM/ML and OPT dominate (cf. Figure 5), and has become increasingly dominant in recent years (cf. Figure 6), yet with several viable alternatives in the shape of PROB, EXP, and SIM. While a pronounced emphasis on tree-based ensembles for predictions and Linear Programming methods for prescriptions was found (cf. Sections 9.1 and 9.2), individual method choices varied widely between applications, exhibiting a textbook level of coverage for the most popular method types. To our knowledge, the various prescriptive EXP methods described in Section 9.4 have not been covered explicitly elsewhere, despite the popularity of prescriptive EXP (cf. Figure 9).
In terms of how methods were combined, we found that DPSA applications pivot around a relatively small number of method combinations among the aforementioned types (cf. Section 10.1), utilizing primarily DM/ML and secondarily PROB for predictions, and primary OPT, supplemented by DM/ML and EXP, for prescriptions. We observe from Figure 9 that there is seemingly a power law at play in terms of popularity of individual method types for prediction/prescription, as well as their combinations, with (combinations of) DM/ML and OPT clearly being the most popular, increasingly over time as well (cf. Figure 10), yet contested by EXP for prescriptions and PROB for predictions.
As for workflow patterns, we identified two generic patterns for structuring DPSA workflows, dividing the set of surveyed papers into two evenly-sized parts (cf. Section 10.2). The Predict-Then-Prescribe pattern is already well-known within the methodological PSA literature in some form [26], as a sequential pattern with a clear-cut separation between prediction and prescription. The alternative Predicting-While-Prescribing pattern, as introduced in this survey, breaks the mold of PTP, bringing prediction and prescription closer together by simultaneous relationships (iterative and/or overlapping), with recent methodological developments even unifying data-driven prediction and automatic prescription in one model.
@@ 11 Synthesis on methods
12 Promising research directions
Based on pervasive patterns found in the set of surveyed papers, we derive four promising research directions (RD1-4) to address RQ3, presented in the following subsections.
12.1 The promise of Big DPSA
Extracting value from large datasets, i.e., Big Data, can be said to form a significant part of a narrative surrounding PSA, in which PSA, extending the reach of its descriptive and predictive counterparts (cf. Figure 1), is the third and final piece to unlock maximum Value from Big Data in a fully automated, large-scale analytics infrastructure [169, 30, 27]. As shown by our temporal analyses in Figures 6 and 10, DM/ML has indeed played an increasingly prominent role within the DPSA field in recent years, both for predictions and prescriptions, and a substantial number of surveyed papers correspondingly discuss the potentials of Big Data in their opening or closing sections [95, 138, 89, 62, 83, 112, 137, 116, 114, 82, 113]. Nonetheless, hardly any DPSA applications in this survey work with what we would even consider to be Big Data.
To clarify our own definition of the phenomenon, we prefer to avoid arbitrary distinctions counted in bytes or cardinalities and instead define data-intensive processing by the need for distributed processing - as exemplified by the primary use cases of distributed data ingestion, analytics and storage systems like Apache Kafka, Spark, and Cassandra [170, 38, 171]. In other words, if a dataset can feasibly be analysed on one workstation/server with respect to application requirements, as done in most surveyed papers (e.g., [87, 109, 126, 129, 78, 84]), then any standard analytics solution like Excel and MySQL might suffice, and said dataset is not considered Big.
A few papers describe their own work as involving Big Data or evidently use distributed processing to handle large datasets [92, 124, 144, 80]. Yet these papers tend to omit implementation details, in some cases possibly for confidential reasons [144], while others only analyse single-node-sized datasets [80].
@@ 12.1 The promise of Big DPSA
12.2 Under-explored application domains
As found in connection to RQ1 (cf. Section 8), DPSA has already found use in a variety of identified application domains, and our temporal analysis demonstrates an ebb and flow in terms of relative prominence among these domains over time (cf. Figure 4). Still, several prominent real-world business domains have been left (almost) unexplored by the set of surveyed papers as a whole. Using UN’s International Standard Industrial Classification (ISIC), revision 4, [172], we have identified eleven business domains in which at most three DPSA applications were found in this survey. Leaving out the miscellaneous sections S, T, and U [172,  p. 43], the eight remaining business domains and section letters, along with associated surveyed papers, if any, are the following:
  • 1.
@@ 12.2 Under-explored application domains
12.3 Methodological alternatives to (mixed-)integer linear programming
While covering OPT methods in Section 9.2, we found that Linear Programming, and in particular (Mixed-)Integer Linear Programming ((M)ILP) algorithms, were the most dominant class of methods by far. What we didn’t dwell on, for the sake of brevity, is the frequently recurring observation that one does not simply use (M)ILP for DPSA. Among DPSA applications there is a strong tendency of various difficulties arising whenever authors try to apply (M)ILP to real-world problems, calling for improvised modelling tricks, heuristics, and workflows.
@@ 12.3 Methodological alternatives to (mixed-)integer linear programming
Another challenge with (M)ILP when considering DPSA specifically, is the popularity of predictive DM/ML (cf. Figure 9), in particular non-parametric or non-linear models (cf. Section 9.1). Limited options for embedding such models into linear optimization models makes (M)ILP less expressive, to the possible detriment of solution quality, e.g., as found in [143].
@@
We recognize that it might be impossible to design a perfect alternative to current prescriptive methods, providing hard guarantees without hard costs. Still, in our temporal analysis of Figure 10 we observe increasing efforts to find good alternative prescriptive methods to rigid OPT strategies piloted by domain experts. Specifically, leveraging DM/ML directly for prescriptions, as manifested in various PTP/PWP workflows (cf. Section 10.2), seems like a promising direction - especially due to how data-driven methods like Deep Learning have already made great strides in solving hard prediction problems while being relatively easy to pilot with high-level tools [39]. There is indeed already a considerable body of work within ML investigating how to approximately solve (especially) combinatorial optimization problems, e.g., Travelling Salesman problems, with neural networks [173]. Yet to our knowledge, none of these works have considered how to incorporate these methods into complete DPSA workflows. Therefore, we propose as a possible research direction to:
RD3: Develop scalable, expressive, usable, and reliable prescriptive DM/ML methods for DPSA applications.
12.4 Flexible tools for DPSA
From our analyses concerning RQ2 (cf. Sections 9 and 10), we note that DPSA applications display a high degree of methodological diversity, both in terms of individual methods and how they are combined, having diversified further over time (cf. Figure 10). Accommodating such a wide variety of workflows poses a challenge to developers of dedicated DPSA tools: As is evident from the aforementioned sections, a PTP pattern with ML for predictions and Linear Programming for prescriptions is but one way to structure a DPSA workflow among many. As for (M)ILP specifically, we furthermore saw in Section 12.3 that significant workflow variations even exist within this one category.
@@ 12.4 Flexible tools for DPSA
13 Conclusion
The primary objective of this survey was to investigate where the many applications of DPSA have been and where they are going (cf. Section 1). In this endeavour, we decided to delimit our scope to Data-Driven PSA, a field deemed to have a sound basis in the historical origins of PSA as well as recent calls for papers (cf. Section 4).
In answering our three research questions in Sections 8-12, we have provided new conceptual models and temporal analyses of the field as a whole, both in terms of problem types (cf. Section 8), method types (cf. Sections 9), and workflow patterns (cf. Section 10). Based on temporal trends and pervasive Gordian Knots found within the set of surveyed papers, we have furthermore provided a foundation for future work by proposing a number of possible research directions (cf. RD1-4 from Section 12), both in terms of new methods, tools, and use cases. We can mnemonically summarize the key contributions in our conceptual framework as 10 applications domains, 5 method types, 2 generic workflow patterns, and 4 research directions. These contributions form a comprehensive model of the current state of the DPSA application research field and where it might be going in the future.
Looking ahead, we believe that this survey can be a useful guide for new recruits and veterans alike. As found in Section 12, there are still ample opportunities for developing new methods, tools, and use cases for DPSA applications with a positive impact on the real world. Revisiting our opening quote in Section 1, we hope that upcoming DPSA applications will continue to find new ways to create the future by predicting it.
Author contributions: CRediT
Martin Moesmann: Conceptualization, Data Curation, Formal Analysis, Investigation, Methodology, Validation, Visualization, Writing – original draft, Writing – review & editing Torben Bach Pedersen: Conceptualization, Funding acquisition, Methodology, Project administration, Supervision, Validation, Writing – review & editing.